a 58-m garden hose is stretched diagonally across a yard. if the width of the yard is 24 m what is the length of the yard round to the nearest tenth
Answer by TimothyLamb(4379)

58 = sqrt( 24*24 + LL )
58*58 = 24*24 + LL
LL = 58*58 - 24*24
LL = 2788
---
L = 52.8 m
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php